Kentucky Section of MAA Annual Meeting Faculty Registration Form

Bellarmine University, Louisville April 4 - 5, 2003

Name: School: Address:

Phone: E-mail: Check all that apply:

Conference Registration and Dues $ 25 Short Course (Friday afternoon) No charge I would like to give a talk at the meeting. No charge Submit the Presentation Abstract Form to Bill Fenton by March 3, 2003. Friday Banquet $ 16 Friday Invited Address No charge Aftermath (Friday evening) No charge Saturday Invited Address No charge Saturday Business Luncheon $ 7

TOTAL ENCLOSED

(checks payable to KY Section of MAA)

Deadline for advance registration is March 20, 2003. On-site registrations are also welcome. Note that the reserved blocks of hotel rooms are released on March 30^{th}. Submit to: Karin Chess, Dept of Math, Owensboro Community College, 4800 New Hartford Road, Owensboro, KY 42303 Karin.Chess@kctcs.edu